OpenStudy (anonymous):

how do you factor y^2-81

OpenStudy (mathstudent55):

Do you recognize the pattern: \(a^2 - b^2 = (a + b)(a - b) \) ?

OpenStudy (anonymous):

yeah

OpenStudy (mathstudent55):

This is called factoring "a difference of trwo squares."

OpenStudy (mathstudent55):

\( y^2 - 81\) \( = y^2 - 9^2\) Now follow the pattern. What do you get?

OpenStudy (anonymous):

y^2-72y-81<-is the answer

OpenStudy (anonymous):

(y+9) (y-9) is the factor?

OpenStudy (mathstudent55):

The second one.

OpenStudy (anonymous):

okay thanks!

OpenStudy (mathstudent55):

wlcm